OpenStudy (anonymous):

find the slope if it exists, on the line y+2.28x-6

OpenStudy (anonymous):

-2.28

OpenStudy (anonymous):

the equation of the line is actually y +2.28x-6=0

OpenStudy (anonymous):

find the slope if it exists, 6x+7y=42

OpenStudy (anonymous):

Take a second to figure out what is going on. You wouldn't have any problems with these. Just solve for y, put y on one side y=mx+b. Once you have that, the slope is the m, the thing in front of x.

OpenStudy (anonymous):

so the slope is 6 ? when I worked the problem it was y=(-6(x-7))/((7))

OpenStudy (anonymous):

You need to work on your algebra a little.\[y =(-6/7)x + 6\]

OpenStudy (anonymous):

so the slope is -6/7 I kinda get it, and yes I do need to work on my algebra, that's why I 'm asking for help

OpenStudy (anonymous):

I'm not putting you down. I am just pointing out why you are having problem with this. Once you get the algebra down this is a breeze.

OpenStudy (anonymous):

okay now Im working on another one the slope if it exists of the line containing the pair of points (-4,-10) and (-5,-11) is this how I do it? (-4,-10)(-5,-11) Subtract 10 from -4, to get -14. (-14)(-5,-11) Subtract 11 from -5, to get -16. (-14)(-16) Multiply -14 by -16 to get 224. (224) 224

OpenStudy (anonymous):

\[m =(y _{2}-y _{1})/(x _{2}-x _{1})\]